Determine the equilibrium levels of all endogenous variables under the assumptions of the classical
macroeconomic framework
1
Expert's answer
2021-08-09T06:48:32-0400

Solution:

The equilibrium levels of all endogenous variables under the classical macroeconomic framework assumptions are determined by solving the model.

The main endogenous variable is income (Y). We first rearrange the model equation such that the endogenous variable is on the left-hand side and the exogenous variables are on the right-hand side, which is called a reduced form.

At equilibrium: Y = AD

Y represents the endogenous variable, while AD represents the exogenous variable

AD = C + I + G + X-M

Therefore, Y = C + I + G + X-M

 

We can also derive equilibrium by equating: Y – C = I + G

Where Y and C are endogenous variables and I and G are exogenous variables